# Unveiling Concept Attribution in Diffusion Models

### v2

@@ Abstract
Diffusion models have shown remarkable abilities in generating realistic and high-quality images from text prompts. However, a trained model remains largely black-box; little do we know about the roles of its components in exhibiting a concept such as objects or styles. Recent works employ causal tracing to localize knowledge-storing layers in generative models without showing how other layers contribute to the target concept. In this work, we approach diffusion models’ interpretability problem from a more general perspective and pose a question: “How do model components work jointly to demonstrate knowledge?”. To answer this question, we decompose diffusion models using component attribution, systematically unveiling the importance of each component (specifically the model parameter) in generating a concept. The proposed framework, called Component Attribution for Diffusion Model (CAD), discovers the localization of concept-inducing (positive) components, while interestingly uncovers another type of components that contribute negatively to generating a concept, which is missing in the previous knowledge localization work. Based on this holistic understanding of diffusion models, we introduce two fast, inference-time model editing algorithms, CAD-Erase and CAD-Amplify; in particular, CAD-Erase enables erasure and CAD-Amplify allows amplification of a generated concept by ablating the positive and negative components, respectively, while retaining knowledge of other concepts. Extensive experimental results validate the significance of both positive and negative components pinpointed by our framework, demonstrating the potential of providing a complete view of interpreting generative models. Our code is available here.
@@
1 Introduction
Recent developments in diffusion models [18, 26, 35, 36] have greatly improved the synthesizing capabilities, including image quality and generating a wide range of knowledge. However, these models lack interpretability; we do not fully understand how they can achieve such impressive performance and how they can generate images from only simple text prompts. To investigate how generative models recall concepts, a recent line of work studies which components in the model store knowledge [27, 2]. In language models, Meng et al. [27] propose causal tracing to locate layers storing facts and reveal that knowledge is localized in middle-layer MLP modules. This method is later transferred to diffusion models in [2], discovering the knowledge distributed hypothesis; this hypothesis shows that, different from language models, knowledge is distributed amongst a set of UNet components and the first self-attention layer of the text-encoder. These approaches shed light on interpreting generative models, enabling more effective model editing  [2, 3]. Nevertheless, they only focus on coarse-grained components (i.e., layers) and knowledge storage – modules that are responsible for generating concepts, potentially ignoring more subtle properties in the generative models and other types of modules, respectively.
This paper first poses a more general question: How do components in diffusion models contribute to a generated concept? We then introduce a framework that predicts the model behavior given the presence of each component based on an efficient linear counterfactual estimator [34]. Through this framework, called Component Attribution for Diffusion Model (CAD), we advance the understanding of how model components activate concepts (e.g., objects, styles, or explicit contents) in diffusion models. In contrast to the prior work that focuses on the model’s layers, CAD allows analysis of more fine-grained components. Specifically, focusing on the most fine-grained components, i.e., the model’s parameters, CAD could also identify concept-inducing (or positive) components similar to knowledge storage; however, instead of the distributed hypothesis in layers [2], CAD discovers the localization hypothesis — knowledge is localized in a small number of parameters. Surprisingly, besides the positive components, CAD also reveals the existences of components that contribute negatively to generating the target concept, which is missing in the previous studies. Ablating these components decreases or increases the probability of generating the corresponding knowledge. This holistic understanding of diffusion models enables a lightweight model editing capability, i.e., to remove (positive) or recall (negative) a concept. Figure 1 illustrates the proposed CAD framework. In summary, our contributions are:
@@ 1 Introduction
  • Utilizing CAD, we confirm the existence of the concept-inducing (positive) model’s parameters, while revealing their localized nature. CAD also uncovers the existence of another type of components – the concept-amplification (negative) components.
  • Leveraging these observations of localized positive and negative components, we develop two lightweight knowledge editing algorithms, CAD-Erase for concept erasing and CAD-Amplify for concept amplification, respectively, for diffusion models.
@@ 2 Related Works
Interpreting Neural Networks. Several research has extensively studied the black-box mechanism of neural networks to explain their behaviors. A line of works [33, 4, 41] aim to visualize important regions in the input of classification models by inspecting the gradient of feature map activations. Sundararajan et al. [37] formalize the problem of attributing the input and propose two axioms to design attribution methods. Fundamentally different from those studies, we seek to attribute the model components, specifically parameters, in diffusion models.
@@
Knowledge Localization. Previous work [16, 8] investigates how components of language model store factual knowledge. Similarly, Shah et al. [34] propose model attribution to analyze the impact of individual components in the image classification and language prediction task, allowing us to understand the relationship between the model components and the model output. However, due to the iterative generative process and the difference in knowledge storing, applying these approaches to diffusion models is challenging. Another line of research [2, 3, 17, 27, 38, 7, 47] utilizes causal analysis to identify critical layers where knowledge is predominantly stored. For instance, Meng et al. [27] show that language models tend to store factual information in certain causal layers, and modifying these layers improves generalization and specificity. Similarly, Basu et al. [2, 3] apply this technique to T2I Latent Diffusion variants, targeting specific text-encoder and U-Net layers to remove unwanted elements such as nudity or copyrighted styles. While these methods have shown successes in localizing knowledge, Hase et al. [17] discover that editing non-causal layers can also modify stored facts in language models. This finding implies that causal analysis may answer a different question from model editing. Furthermore, these approaches inspect the activations, which are dependent on the input, whereas our work studies the parameters of the model. Another study [9] examines the weight space of customized diffusion models by creating a dataset of thousands of models, showing it behaves as an interpretable manifold for generating and modifying visual identities. In contrast, our work offers an efficient approach to studying the role of the model component.
@@
Concept Erasure. Due to the reliance on large, uncontrollable datasets, latent diffusion models (LDMs) may generate undesirable content, such as nudity, outdated information, or copyrighted artistic styles. Early effort address this problem by fine-tuning Cross-Attention layers to minimize the unlearn losses [11, 20, 21, 48, 29], or editing the text-encoder  [1, 2] to remove unwanted concepts. In addition, several studies [12, 25, 42] point out the necessity of removing multiple concepts simultaneously in real-world scenarios. Specifically, Gandikota et al. [12] and Lu et al. [25] extend cross-attention fine-tuning via a closed-form solution, while Xiong et al. [42] target MLP layers in the text-encoder. Recent studies aim to improve robustness of erasing methods to red-teaming attacks, including ConceptPrune [5], RECE [13], RACE [19], and pruning methods [43]. These methods enable efficient erasure of various contents while ensuring minimal interference with the unedited ones.
@@
Concept Amplification. Motivated by Dreambooth [32], Cones [24] inserts new objects into the model by examining the gradient to find concept neurons. Different from this work, CAD-Amplify locates components to magnify existing knowledge in diffusion models. Dai et al. [8] also propose a method to amplify facts, but relies on amplifying positive neurons. Our work is the first study showing the existence of negative components and how to systematically locate them.
@@
Red-Teaming Attacks. Although model fine-tuning eliminates undesirable concepts in text-to-image models, recent studies [45, 6, 49, 44, 46, 40, 30] demonstrate that this approach remains unreliable against adversarial prompt attacks. Black-box attacks such as SneakyPrompt [45], Ring-A-bell [40], and MMA-Diffusion [44] can bypass many existing safety mechanisms without assuming access to the model’s parameters, by creating unsafe prompts with similar embeddings. Concurrently, several white-box attacks such as P4D [6], UnlearnDiff [49], and CCE [30] can also induce the fine-tuned models to regenerate sensitive outputs, by crafting adversarial prompts. These attacks highlight the need for robust defenses, which can completely remove undesirable concepts and preserve the quality of generated images. More importantly, we can also employ these attacks to examine the existence of a concept in a diffusion model, enabling assessment of whether a method can fully erase this concept.
@@
Pruning Approaches. Similar to our algorithms, many studies have investigated the effect of pruning neural networks, however, for the time and memory efficiency [14, 10]. Specifically, [28, 22, 39] utilize the gradient information to measure the parameters’ importance and remove unnecessary parameters with low scores to improve inference speed. In contrast, we propose two algorithms removing parameters that have the most significant positive or negative contributions to erase or amplify knowledge.
@@ 4.1 Decomposing Knowledge in Diffusion
Our goal is to interpret how each component $w_{i}$ contributes to generating a concept, quantified by $J(c,w)$. Specifically, we estimate how $J(c,w)$ changes if we remove a component $w_{i}$, i.e. setting its value to 00. Let $\tilde{w}$ be the new set of components obtained by adjusting some components to 00, we want to find a function $g(\boldsymbol{0}_{\tilde{w}};c)=J(c,\tilde{w})$ where $\boldsymbol{0}_{\tilde{w}}\in\{0,1\}^{d}$, $d$ is the number of components, and
@@
Each coefficient $\alpha_{c,i}$ represents how the component $w_{i}$ contributes to the concept $c$.
@@ 4.2 CAD: Component Attribution for Diffusion Model
$\smash{\sum_{i\in S}\alpha_{c,i}}$ $=J(c,w)-J(c,\tilde{w})$
$\approx(w-\tilde{w})\nabla_{w}J(c,w)=\sum_{i\in S}w_{i}\frac{\partial J(c,w)}{\partial w_{i}}.$ (3)
@@
Algorithm 1 Erasing knowledge in generative models
@@
Compute attribution scores $w_{i}\frac{\partial J}{\partial w_{i}}$ with Equation (4).
Locate top-$k$ components $w_{i}\in S$ with the highest (positive) attribution.
@@
Algorithm 2 Amplifying knowledge in generative models
Diffusion model $\Phi$, target concept $c$, the number of components $k$, the set of images $x$ of concept $c$.
Compute attribution scores $w_{i}\frac{\partial J}{\partial w_{i}}$ with Equation (5).
Set $w_{i}\leftarrow 0,w_{i}\in S$
5 Editing Diffusion Models with CAD
In this section, we investigate the application of CAD and study how the model parameters impact concept generation. We then also propose two lightweight, inference-time editing algorithms that remove (CAD-Erase) or amplify (CAD-Amplify) a concept in diffusion models.
5.1 Localizing and Erasing Knowledge
@@
5.2 Amplifying Knowledge in Diffusion Models
Hypothesis 2.
Negative components exist and ablating them will amplify knowledge.
Figure 2: The attribution scores predicted by CAD and the actual values of the objective function.
6 Experiments
In this section, we aim to verify and provide a comprehensive empirical analysis of the knowledge localization hypothesis in Section 6.2 and the existence of negative components in Section 6.3.
6.1 CAD Well Approximates the Change in the Objective
First, we evaluate how good the proposed first-order approximation is and whether CAD can accurately capture component attributions. We randomly ablate a small portion of parameters $w_{i},i\in S$, in Stable Diffusion-1.4 and obtain the corresponding change in the objective. We also use CAD to compute the predicted change, indicated by $\sum_{i\in S}w_{i}\frac{\partial J}{\partial w_{i}}$. We repeat this process $1000$ times and evaluate CAD. Figure 2 illustrates that our predicted values estimate well the actual changes in the objective with a high Pearson correlation. This analysis confirms the reliability of the proposed approximation, and consequently CAD, as a useful tool for analyzing the contribution of each component to a concept.
@@ 6.2 CAD Can Locate Positive Components and Erase Knowledge
The analysis in the previous section shows that CAD can successfully identify positive and negative components. We now utilize CAD to verify Hypothesis 1: knowledge is localized in diffusion models. We conduct experiments on Stable Diffusion-1.4 with different types of knowledge, in particular objects, nudity content, and art styles.
@@
Figure 3: The qualitative results of CAD. Removing positive components to “English Springer” prevents diffusion from generating that concept. Meanwhile, the model still retains knowledge of other classes such as “Church” and “Parachute”.
Figure 4: The first two rows contain images generated by the original model and erasing methods on I2P prompts. We add * for publication. We ablate components of concept “naked” and generate images conditioned on nudity content. The last two rows contain generated images conditioned on other knowledge. Prompt (1): “High-resolution photograph of a bustling street market, vibrant colors”; Prompt (2): “Realistic landscape of rolling hills, vibrant greenery”.
@@
Table 3: The number of nudity content and the drop in percentage from the original model classified by Nudenet 3.4.2 on images generated from adversarial prompts. Lower is better.
Model MMA Ring-a-bell
SD-1.4 1941 (--00.00%) 414 (--00.00%)
ConceptPrune 98 (--94.95%) 83 (--79.95%)
ESD 279 (--85.62%) 95 (--77.05%)
RECE 481 (--75.22%) 4 (--99.03%)
UCE 971 (--49.97%) 64 (--84.54%)
CAD-Erase 62 (--96.81%) 5 (--98.79%)
@@
Artist LPIPS on the target $\uparrow$ LPIPS on other $\downarrow$
ESD RECE UCE CAD-Erase ESD RECE UCE CAD-Erase
@@
Erasing with adversarial prompts. Recent works [45, 40, 44] show that current erasing methods do not completely remove knowledge from the model, and propose attack methods that create adversarial prompts to induce the erased model to still generate harmful content. We evaluate our method on two unsafe prompt sets, MMA and Ring-A-Bell, in Table 3. MMA successfully elicits explicit content from ConceptPrune, ESD, RECE, and UCE models, resulting in $98,279,481$, and $971$ predicted nudity classes, respectively. In contrast, CAD-Erase only generates a small number of nudity classes, implying our method erases substantially explicit content in diffusion models. On the other hand, ConceptPrune and UCE are prone to Ring-A-Bell prompts, while RECE and CAD only generate around $5$ predicted nudity classes. We also evaluate the model with white-box attacks [6, 15]. Table 4 reports the attack success rate of white-box attacks in making the erased model generate the target concept. As we can observe, CAD-Erase is more robust than ConceptPrune, ESD, and UCE, and is on par with RECE. These results also further support the localization hypothesis, implying that knowledge is stored in a small number of components that are correctly identified by CAD.
@@
Erasing art styles. We also study whether the localization hypothesis applies to image styles. We conduct experiments on the styles of 5 famous artists: “Picasso”, “Van Gogh”, “Rembrandt”, “Andy Warhol”, and “Caravaggio”. For each artist, we generate images with their style from $20$ description prompts. We report the LPIPS score of images generated by SD-1.4 and the model created by CAD and other erasing methods in Table 5. Figure 5 illustrates qualitative results of CAD on the target artist and other artists. Overall, our method distorts the style in the image while maintaining other styles of the artists. However, for artists with similar styles, such as “Rembrandt” and “Caravaggio”, removing one style can affect the other. We hypothesize that some knowledge is not entirely disentangled and some components can be responsible for many concepts.
Table 7: The number of nudity content detected by Nudenet, generated by models in which nudity is amplified by CAD.
Model Armpits Belly Buttocks Feet Breast (F) Genitalia (F) Breast (M) Genitalia (M) Anus Total
SD-1.4 169 197 26 28 271 29 60 18 0 798
CAD-Amplify 229 242 31 34 360 33 44 18 0 991
ESD 17 15 6 4 22 12 1 11 0 88
CAD-Amplify (ESD) 26 19 8 3 19 11 2 14 0 102

### v3

@@ Abstract
Diffusion models have shown remarkable abilities in generating realistic and high-quality images from text prompts. However, a trained model remains largely black-box; little do we know about the roles of its components in exhibiting a concept such as objects or styles. Recent works employ causal tracing to localize knowledge-storing layers in generative models without showing how other layers contribute to the target concept. In this work, we approach diffusion models’ interpretability problem from a more general perspective and pose a question: “How do model components work jointly to demonstrate knowledge?”. To answer this question, we decompose diffusion models using component attribution, systematically unveiling the importance of each component (specifically the model parameter) in generating a concept. The proposed framework, called Component Attribution for Diffusion Model (CAD), discovers the localization of concept-inducing (positive) components, while interestingly uncovers another type of components that contribute negatively to generating a concept, which is missing in the previous knowledge localization work. Based on this holistic understanding of diffusion models, we present and empirically evaluate one utility of component attribution in controlling the generation process. Specifically, we introduce two fast, inference-time model editing algorithms, CAD-Erase and CAD-Amplify; in particular, CAD-Erase enables erasure and CAD-Amplify allows amplification of a generated concept by ablating the positive and negative components, respectively, while retaining knowledge of other concepts. Extensive experimental results validate the significance of both positive and negative components pinpointed by our framework, demonstrating the potential of providing a complete view of interpreting generative models. Our code is available here.
1 Introduction
Recent developments in diffusion models [18, 26, 36, 37] have greatly improved the synthesizing capabilities, in terms of image quality and the diversity of generated knowledge. However, these models lack interpretability; we do not fully understand how they can translate simple prompts to coherent visual outputs. To investigate how generative models express learned concepts, a recent line of work studies which components in the model store knowledge [2, 27]. In language models, Meng et al. [27] propose causal tracing to locate layers storing facts and reveal that knowledge is localized in middle-layer MLP modules. Basu et al. [2] transfer this approach to diffusion models and propose the knowledge distributed hypothesis. They show that, unlike language models, knowledge is distributed among a set of UNet components and the first self-attention layer of the text encoder. These works shed light on interpreting generative models, enabling more effective model editing  [3, 2]. Nevertheless, they focus only on knowledge storage—modules that are responsible for generating concepts—and on coarse-grained components, such as layers. This perspective may overlook more subtle properties and other types of modules that also influence the outputs.
@@ 1 Introduction
To address that limitation, this paper first poses a more general question: How do components in diffusion models contribute to a generated concept? We introduce a framework that predicts the model’s behavior given the presence of each component by identifying its contribution through an efficient linear counterfactual estimator [35]. Through this framework, called Component Attribution for Diffusion Model (CAD), we advance the understanding of how model components activate concepts (e.g., objects, styles, or explicit contents) in diffusion models. In contrast to the prior work that focuses on the model’s layers, CAD allows analysis of more fine-grained components. Specifically, we revisit the distributed hypothesis [2] and leverage CAD  to identify concept-inducing (or positive) components that are similar to knowledge storage. However, by focusing on the most fine-grained components, i.e., the model’s parameters, we propose the localization hypothesis: knowledge is localized in a small number of parameters. Surprisingly, besides the positive components, CAD also reveals the existence of components that contribute negatively to generating the target concept, which is missing in the previous studies. Ablating positive or negative components decreases or increases the probability of generating the corresponding knowledge. As one example of its utility, this holistic understanding of diffusion models enables a lightweight model editing capability, i.e., to remove (positive) or recall (negative) a concept. Figure 1 illustrates the proposed CAD framework. In summary, our contributions are:
@@
  • Utilizing CAD, we confirm the existence of the concept-inducing (positive) model’s parameters, while revealing their localized nature. CAD is also the first work that uncovers the existence of another type of components – the concept-amplification (negative) components.
  • Leveraging these observations of localized positive and negative components, we develop two lightweight knowledge editing algorithms for diffusion models, CAD-Erase for concept erasing and CAD-Amplify for concept amplification, respectively.
@@ 2 Related Works
Interpreting Neural Networks. Several research has extensively studied the black-box mechanism of neural networks to explain their behaviors. A line of works [34, 4, 42] visualize important input regions of classification models by using the gradient of feature map activations. Sundararajan et al. [38] formalize the problem of attributing the input and propose two axioms to design attribution methods. Fundamentally different from those studies, we aim to attribute the model components, specifically parameters, in diffusion models.
@@
Knowledge Localization. Previous work explored how language model components store factual knowledge [16, 8] or used model attribution to analyze the impact of individual components in the image classification and language prediction task[35]. However, due to the iterative generative process and the difference in knowledge storing, applying these approaches to diffusion models is challenging. Another line of research [2, 3, 17, 27, 39, 7, 49] utilizes causal analysis to identify critical layers for knowledge in language models and T2I Latent Diffusion variants. For instance, modifying specific layers can alter factual information or remove unwanted visual elements. While these methods have shown successes in localizing knowledge, Hase et al. [17] discover that editing non-causal layers can also modify stored facts in language models. This finding implies that causal analysis may answer a different question from model editing. Furthermore, these approaches inspect the activations, which are dependent on the input, whereas our work studies the parameters of the model. Dravid et al. [9] examine the weight space of several customized diffusion models; in contrast, our work offers an efficient approach to studying individual model component roles.
@@
Concept Erasure. Latent diffusion models (LDMs) can generate undesirable content (e.g., nudity, outdated information, copyrighted artistic styles) due to their large and uncontrolled training datasets. Early efforts address this problem involved fine-tuning Cross-Attention layers  [11, 20, 21, 50, 29] or editing the text-encoder  [1, 2]. In addition, several research [12, 25, 44] highlight the necessity to remove multiple concepts simultaneously in real-world scenarios. More recent works aim to improve robustness of erasing methods to red-teaming attacks, including ConceptPrune [5], RECE [13], RACE [19], and pruning methods [45]. These methods enable efficient erasure of various contents while ensuring minimal interference with the unedited ones.
@@
Concept Amplification. Motivated by Dreambooth [33], Cones [24] inserts new objects into the model by identifying concept neurons. In contrast, CAD-Amplify locates components to magnify existing knowledge in diffusion models. Dai et al. [8] also proposes a method to amplify facts, but relies on amplifying positive neurons. Our work is the first study showing the existence of negative components and how to systematically locate them.
@@
Red-Teaming Attacks. Although fine-tuning eliminates undesirable concepts in text-to-image models, recent studies [47, 6, 51, 46, 48, 41, 30] show that this approach remains unreliable against adversarial prompt attacks. These safety mechanisms can be bypassed by both black-box (e.g., SneakyPrompt [47], Ring-A-bell [41]) and white-box attacks (e.g., P4D [6], UnlearnDiff [51]), leading to the regeneration of sensitive content. These attacks highlight the need for robust defenses that fully remove concepts while preserving image quality. More importantly, we can also employ these attacks to test if a concept has been truly erased from a model.
@@
Pruning Approaches. Similar to our algorithms, many studies  [14, 10] have investigated pruning neural networks, primarily for time and memory efficiency. Specifically, [28, 22, 40] use gradient information to identify and remove less important parameters, thereby improving inference speed. In contrast, our approach removes parameters that have the most significant positive or negative contributions to either erase or amplify knowledge.
@@ 4.1 Decomposing Knowledge in Diffusion
[…]
@@
Each coefficient $\alpha_{c,i}$ represents how the component $w_{i}$ contributes to the concept $c$. Another way to intervene is to amplify the effect of $w_{i}$ by rescaling its magnitude; however, unlike knockout, which removes the component completely, scaling may preserve interactions, and deciding the magnitude of the change can be non-trivial. Nevertheless, we investigate this type of intervention in the Appendix.
@@ 4.2 CAD: Component Attribution for Diffusion Model
$\smash{\sum_{i\in S}\alpha_{c,i}}=J(c,w)-J(c,\tilde{w})\approx(w-\tilde{w})\nabla_{w}J(c,w)=\sum_{i\in S}w_{i}\frac{\partial J(c,w)}{\partial w_{i}}.$ (3)
@@
5 Editing Diffusion Models with CAD
In this section, we discuss the importance of studying parameter attributions in concept generation via two applications of CAD. Specifically, we propose and empirically evaluate two lightweight, inference-time editing algorithms that remove (CAD-Erase) or amplify (CAD-Amplify) a concept in diffusion models.
5.1 Localizing and Erasing Knowledge
@@
Compute the scores $w_{i}\frac{\partial J}{\partial w_{i}}$ with Eq. (4).
Locate top-$k$ components $w_{i}\in S$ with the (positive) attribution.
@@
Algorithm 1 CAD-Erase
@@
Diffusion model $\Phi$, target concept $c$, the n.o. components $k$, images $x$ of concept $c$.
Compute the scores $w_{i}\frac{\partial J}{\partial w_{i}}$ with Eq. (5).
Set $w_{i}\leftarrow 0,w_{i}\in S$
Algorithm 2 CAD-Amplify
@@
5.2 Amplifying Knowledge in Diffusion Models
Hypothesis 2.
Negative components exist, and ablating them will amplify knowledge.
Figure 2: The attribution scores predicted by CAD and the actual values of the objective.
6 Experiments
In this section, we aim to verify and provide a comprehensive empirical analysis of the knowledge localization hypothesis in Section 6.2 and the existence of negative components in Section 6.3. We provide additional results on other diffusion models, different intervention, and ablating on different modules in the Appendix.
6.1 CAD Well Approximates the Change in the Objective
@@
In diffusion models, as mentioned in Section 4, attributing the components is time-consuming and more complicated due to their iterative generation process. Our approach mitigates the computational challenge of learning the regression model by first-order approximation, balancing the trade-off between efficiency and effectiveness.
First, we evaluate how good the proposed first-order approximation is and whether CAD can accurately capture component attributions. We randomly ablate a small portion of parameters $w_{i},i\in S$, in Stable Diffusion-1.4 and obtain the corresponding change in the objective. We also use CAD to compute the predicted change, indicated by $\sum_{i\in S}w_{i}\frac{\partial J}{\partial w_{i}}$. We repeat this process $1000$ times and evaluate CAD. Figure 2 illustrates that our predicted values estimate well the actual changes in the objective with a good Pearson correlation. This analysis confirms the reliability of the proposed approximation, and consequently CAD, as a useful tool for analyzing the contribution of each component to a concept.
@@ 6.2 CAD Can Locate Positive Components and Erase Knowledge
The analysis in the previous section shows that CAD can successfully identify positive and negative components. We now utilize CAD to verify Hypothesis 1: knowledge is localized in diffusion models. We conduct experiments on Stable Diffusion-1.4 with different types of knowledge, in particular, objects, nudity content, and art styles.
@@
Figure 3: The qualitative results of CAD. Removing positive components to “English Springer” avoids generating that concept. Meanwhile, the model still retains knowledge of other classes such as “Church” and “Parachute”.
Figure 4: The first two rows contain images generated by the original model and erasing methods on I2P prompts. We add * for publication. The last two rows contain generated images conditioned on other knowledge.
@@
CAD-Amplify 229 242 31 34 360 33 44 18 0 991
@@
Table 4: The number of nudity content and the drop from the original model classified by Nudenet on images generated from adversarial prompts. Lower is better.
Model MMA Ring-a-bell
SD-1.4 1941 (-00.00%) 414 (-00.00%)
ConceptPrune 98 (-94.95%) 83 (-79.95%)
ESD 279 (-85.62%) 95 (-77.05%)
RECE 481 (-75.22%) 4 (-99.03%)
UCE 971 (-49.97%) 64 (-84.54%)
CAD-Erase 62 (-96.81%) 5 (-98.79%)
Erasing with adversarial prompts. Recent works [47, 41, 46] show that current erasing methods do not completely remove knowledge from the model, and propose attack methods that create adversarial prompts to induce the erased model to still generate harmful content. We evaluate our method on two unsafe prompt sets, MMA and Ring-A-Bell, in Table 6.2. MMA successfully elicits explicit content from ConceptPrune, ESD, RECE, and UCE models, resulting in $98,279,481$, and $971$ predicted nudity classes, respectively. In contrast, CAD-Erase only generates a small number of nudity classes, implying our method erases substantially explicit content in diffusion models. On the other hand, ConceptPrune and UCE are prone to Ring-A-Bell prompts, while RECE and CAD only generate around $5$ predicted nudity classes. We also evaluate the model with white-box attacks such as P4D [6] and UnlearnDiff (UD) [15]. Table 6.2 reports the attack success rate of white-box attacks in making the erased model generate the target concept. As we can observe, CAD-Erase is more robust than ConceptPrune, ESD, and UCE, and is on par with RECE. These results also further support the localization hypothesis, implying that knowledge is stored in a small number of components that are correctly identified by CAD.
Erasing art styles. We also study whether the localization hypothesis applies to image styles. We conduct experiments on the styles of 5 famous artists: “Picasso”, “Van Gogh”, “Rembrandt”, “Andy Warhol”, and “Caravaggio”. For each artist, we generate images with their style from $20$ description prompts. We report the LPIPS score of images generated by SD-1.4 and the model created by CAD and other erasing methods in Table 6.2. Figure 5 illustrates qualitative results of CAD on the target artist and other artists. Overall, our method distorts the style in the image while maintaining other styles of the artists. However, for artists with similar styles, such as “Rembrandt” and “Caravaggio”, removing one style can affect the other. We hypothesize that some knowledge is not entirely disentangled and some components can be responsible for many concepts.
@@
[…]
ESD RECE UCE CAD ESD RECE UCE CAD